OpenStudy (anonymous):

Looking for help evaluating limits with L'Hôpital's Rule Hello. For an exercise, I'm supposed to evaluate \[ \lim_{x \to \infty} \frac{(\ln \ x)^2}{x^3} \] using L'Hôpital's Rule. I'm having trouble finding the answer since I haven't worked with limits since the beginning of my calc course. So far I have: (continued in next post...)

OpenStudy (anonymous):

So far I have: \[\begin{align*} \lim_{x \to \infty} \frac{ (\ln\ {x})^2 }{x^3} &= \lim_{x \to \infty} \frac{ \frac{d}{dx} \left[ (\ln\ {x})^2 \right] }{ \frac{d}{dx} \left [x^3 \right]}\\ \\ &= \lim_{x \to \infty} \frac{ \frac{2 \ln(x)}{x} }{3 x^2}\\ \\ &= \lim_{x \to \infty} \frac{ \frac{2}{x^2} - \frac{2 \ln(x)}{x^2} }{ 6x }\\ \\ &= \lim_{x \to \infty} \frac{ \frac{4 \ln(x) - 6}{x^3}}{6} \end{align*}\] It doesn't look like just applying L'Hôpital's rule works; does the function need to be rewritten or something? Any help is appreciated; thanks in advance!

myininaya (myininaya):

have you proved ln(x)/x goes to inf as x goes to inf ?

myininaya (myininaya):

that actually goes to 0 not inf so on that one step after the first round of l'hospital you have 0/inf

myininaya (myininaya):

you can play with it though

myininaya (myininaya):

to get it as inf/inf

OpenStudy (kainui):

I am not sure what you did between these two steps here: \[\begin{align*} \\ &= \lim_{x \to \infty} \frac{ \frac{2 \ln(x)}{x} }{3 x^2}\\ \\ &= \lim_{x \to \infty} \frac{ \frac{2}{x^2} - \frac{2 \ln(x)}{x^2} }{ 6x }\\ \end{align*}\] It looks like you should just combine them to get: \[\begin{align*} \\ &= \lim_{x \to \infty} \frac{ 2 \ln(x) }{3 x^3}\\ \end{align*}\]

myininaya (myininaya):

that is the form I'm speaking of ^ the inf/inf thingy! :)

OpenStudy (anonymous):

myininaya: Nope, I haven't proven that. The limit of \(\frac{\ln x}{x}\) as x goes to infinity is 0, correct? Maybe that's where I messed up. But I don't understand why that limit is 0; it's not in my book

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{\ln(x)}{x}=\lim_{x \rightarrow \infty}=\frac{\frac{1}{x}}{1}=\lim_{x \rightarrow \infty}\frac{1}{x}=0 \\ \text{ so you can't use l'hospital on this step } \\ \lim_{x \rightarrow \infty} \frac{\frac{2\ln(x)}{x}}{3x^2}\] until you rewrite it as the form @kainui did

myininaya (myininaya):

since you would have 0/inf

OpenStudy (anonymous):

Kainui: Instead of combining the numerator and denominator in \[ \lim_{x \to \infty} \frac{ \frac{2 \ln(x)}{x} }{3 x^2}, \] I derived the numerator and denominator without simplifying.

OpenStudy (anonymous):

My book isn't very clear on when to use l'hopital; is the function supposed to have the form \[ \frac{\infty}{\infty}? \]

myininaya (myininaya):

yea that is one form

myininaya (myininaya):

\[\pm \frac{\infty}{\infty}, \frac{0}{0}\]

OpenStudy (anonymous):

Hmm okay. What about forms like \( \frac{\infty}{0} \text{or} \frac{0}{\infty}?\) Are those indeterminate forms that you can use l'hopital on?

myininaya (myininaya):

no you can't use l'hospital on those unless you can rewrite it in the forms I mentioned

myininaya (myininaya):

and then after you have rewritten them in the forms I have mentioned then you can apply l'hospital

OpenStudy (anonymous):

Okay, I think I understand. I'll try Kainui's rewrite

myininaya (myininaya):

cool stuff

OpenStudy (anonymous):

\[ \begin{align*} \lim_{x \to \infty} \frac{ \frac{2 \ln(x)}{x} }{3x^2} &= \lim_{x \to \infty} \frac{2 \ln(x)}{3x^3}\\ \\ &= \lim_{x \to \infty} \frac{ \frac{2}{x} }{9x^2}\\ \\ &= \lim_{x \to \infty} \frac{2}{9x^3} \end{align*}\] Okay, so once it's \( \frac{2}{9x^3} \) that means I cannot do l'hopital anymore right?

myininaya (myininaya):

right

myininaya (myininaya):

and it also means you are done

myininaya (myininaya):

because 1/inf->

myininaya (myininaya):

question mark comes after that :p

myininaya (myininaya):

1/inf->? there that sounds more complete :p

OpenStudy (anonymous):

1 / inf = 0 correct?

myininaya (myininaya):

yep yep

myininaya (myininaya):

so any constant times that will go to that constant *0 which is just 0

OpenStudy (anonymous):

Ohh... I think I understand. I do have one more (quick) question if you're willing

myininaya (myininaya):

\[\frac{2}{9}\lim_{x \rightarrow \infty}\frac{1}{x^3}=\frac{2}{9} \cdot 0=0\]

myininaya (myininaya):

k

myininaya (myininaya):

I think I have time for one more

OpenStudy (anonymous):

Thanks :) I was given \[ \lim_{x \to 0^+} \left( \frac{10}{x} - \frac{3}{x^2} \right) \] which can be rewritten as \[ \lim_{x \to 0^+} \left( \frac{10x - 3}{x^2}\right). \] They were saying you could solve that using direct substitution, but that gives you \[ \frac{0 - 3}{0} \] and I thought that was indeterminate

myininaya (myininaya):

you would -3/0 which means it goes to -inf or inf if x is approaching 0 from the right that means x>0 so x^2>0 and 10x>0 10x-3>-3 10x-3>-3 means 10x-3 is near negative numbers as we approach 0 from the right and we know x^2 is always positive

myininaya (myininaya):

like basically you have a neg/pos and you already know it goes to one of the infinities

OpenStudy (anonymous):

Ahh okay. I was trying to think of it strictly analytically instead of graphically. Thanks myininaya :) I think I understand now

myininaya (myininaya):

so for that one you know that neg/pos=neg right?

OpenStudy (anonymous):

Right. Cause if x is approaching from the right, it's positive, and I already have a negative in the numerator, so neg / pos would be a negative value

myininaya (myininaya):

yep yep also the only factor that determines if it positive or negative is the top since we know the bottom is always positive so we only need to look at 10x-3 as x approaches 0 from the right 10(0)-3=-3 So 10x-3 is negative for either direction (right or left)

myininaya (myininaya):

(right or left of 0)*

myininaya (myininaya):

\[\lim_{x \rightarrow 0^-}\frac{|x|}{x}\] this one is a little trickier

myininaya (myininaya):

if x is approaching 0 from the left then x<0 so |x|=-x since x<0

myininaya (myininaya):

\[\lim_{x \rightarrow 0^-}\frac{|x|}{x}=\lim_{x \rightarrow 0^-}\frac{-x}{x}=-\lim_{x \rightarrow 0}\frac{x}{x}=-(1)=-1\]

OpenStudy (anonymous):

I got ya :) that method of solving the limit of abs(x) / x would be easier than l'hopital in my opinion

myininaya (myininaya):

one more example for fun \[\lim_{x \rightarrow 1^+}\frac{x^2-2x-3}{x-1} \\ \lim_{x \rightarrow 1^+}\frac{(x-3)(x+1)}{x-1}\] this means we want to look at x>1 since we are approaching 1 from the right subtract one on both sides x-1>0 so the bottom is positive as we approach 1 from right since x-1>0 means x-1 is positive x-1>0 subtract 2 on both sides x-3>-2 x-3>-2 well this means x-3 is near -2 as we approach 0 from the right so we can look at this factor is if it were negative and now if x>1 and we add 1 on both sides then we have x+1>2 x+1 is near positive 2 when we approach 1 from the right so we already know the limit is one of the infinities now let's look at which one: \[\frac{ \text{ neg }\cdot \text{ pos}}{\ \text{ pos} } \infty\]

myininaya (myininaya):

-2 as we approach 1 from the right*

myininaya (myininaya):

that is how I figure out the sign anyway

myininaya (myininaya):

some other person might have a easier way

OpenStudy (anonymous):

Easier to understand than my textbook, haha. I appreciate you helpin' out and giving other examples :)

myininaya (myininaya):

np

myininaya (myininaya):

there is also the other way like the last one I gave we knew it was one of the infinities and we knew x approaches 1 from the right you can use the cheating way (i call it the cheating way) the pluggin in some value close to 1 from the right like say 1.5 and see if the sign of the output is negative or positive

OpenStudy (rational):

initially you may use graphs to convince more about why the limit is whatever it is..

OpenStudy (anonymous):

Yeah, I'll do that sometimes if I'm confused on a part of the limit.

OpenStudy (rational):

if you graph that function you will see that something bad happens to the graph at x=1

myininaya (myininaya):

vertical asymptotes are bad :p

myininaya (myininaya):

I guess I would consider continuous functions of good behavior :p

OpenStudy (anonymous):

Ah you mean for (x^2 - 2x - 3) / (x-1) ?

myininaya (myininaya):

yeah

myininaya (myininaya):

whenever you have f(a) gives you c/0 where c is not 0 you have a vertical asymptote at x=a

OpenStudy (anonymous):

Okay, but can't you just find the vertical asymptote by setting the denominator equal to 0 and solving? I don't think you need to graph it

OpenStudy (rational):